# A Survey on Self-play Methods in Reinforcement Learning

### v3

@@ Abstract
Self-play, characterized by agents’ interactions with copies or past versions of themselves, has recently gained prominence in reinforcement learning (RL). This paper first clarifies the preliminaries of self-play, including the multi-agent reinforcement learning framework and basic game theory concepts. Then, it provides a unified framework and classifies existing self-play algorithms within this framework. Moreover, the paper bridges the gap between the algorithms and their practical implications by illustrating the role of self-play in different scenarios. Finally, the survey highlights open challenges and future research directions in self-play. This paper is an essential guide map for understanding the multifaceted landscape of self-play in RL.
11footnotetext: Corresponding authors.
1 Introduction
@@
Reinforcement learning (RL) represents a significant paradigm (Sutton & Barto, 2018) within machine learning (ML), focused on optimizing decision processes through interaction with an environment. In RL, the environment is modeled as a Markov decision process (MDP), where agents observe states, take actions, receive rewards, and cause transitions to new states. The primary goal of RL algorithms is to derive the optimal policy that yields the maximum expected accumulated reward over time. Deep RL extends traditional RL by employing deep neural networks as function approximators to handle high-dimensional state spaces, contributing to breakthroughs in various complex tasks (Mnih et al., 2013).
Moreover, transitioning from single-agent to multi-agent reinforcement learning (MARL) introduces complex dynamics (Rashid et al., 2018; Mahajan et al., 2019; Yu et al., 2022). In MARL, the interdependence of agents’ actions presents significant challenges, as the environment appears non-stationary to each agent. The main issues in MARL are coordination, communication, and equilibrium selection, particularly in competitive scenarios. These challenges often lead to difficulties in achieving convergence, maintaining stability, and efficiently exploring the solution space.
@@
With the help of game theory, a mathematical framework that models the interactions between multiple decision-makers, self-play emerges as an elegant solution to some inherent challenges in MARL. Self-play provides an approach where an agent interacts with copies or past versions of itself (Samuel, 2000; Bansal et al., 2018). This method promises a more stable and manageable learning process. The capabilities of self-play extend to a wide range of scenarios, including its high-profile applications in Go (Silver et al., 2016; 2017; 2018; Schrittwieser et al., 2020), chess (Silver et al., 2018; Schrittwieser et al., 2020), poker (Moravčík et al., 2017; Heinrich et al., 2015), and video games (Berner et al., 2019; Vinyals et al., 2019). In these scenarios, it has developed strategies that surpass human expertise. Although the application of self-play is extensive and promising, it is accompanied by limitations, such as the potential convergence to suboptimal strategies and significant computational requirements(Silver et al., 2016; 2018).
@@
Although some research adopts a broad perspective through empirical game-theoretic analysis (EGTA)(Wellman et al., 2025), relatively few comprehensive surveys focus solely on self-play algorithms. Among them, a study develops an algorithmic framework for self-play, but it does not incorporate the Policy-Space Response Oracle (PSRO) series of algorithms (Hernandez et al., 2021), while a separate study focuses exclusively on PSRO (Bighashdel et al., 2024), without considering other self-play algorithms. Another study examines the theoretical safety of self-play (DiGiovanni & Zell, 2021). While these studies are valuable, they do not provide a comprehensive perspective that fully captures the breadth and depth of self-play. Therefore, this survey aims to fill this gap.
@@
The survey is organized as follows. Sec. 2 introduces the background of self-play, including the RL framework and game theory concepts. Sec. 3 proposes a unified framework and then categorizes existing self-play algorithms into four categories based on this framework. In Sec. 4, a comprehensive analysis illustrates how self-play is applied in various scenarios. Sec. 5 describes open problems in self-play and explores future research directions. Finally, Sec. 6 concludes the survey on self-play. A more detailed overview of our survey is depicted in Fig. 1, which further illustrates the relationships among the subsections.
Figure 1: Overview of our survey.
2 Preliminaries
@@
This section first introduces the framework of RL, where an agent learns to make decisions by interacting with an environment to maximize cumulative rewards. We then present basic game theory concepts, which analyze strategic interactions between rational decision-makers and provide tools to study outcomes based on players’ strategies.
2.1 RL Framework
@@
In RL, the environment is modeled as an MDP, where the Markovian assumption states that the environment’s evolution is fully determined by its current state, eliminating the need to consider past states. An agent interacts with the environment by taking actions, which result in different states and associated rewards. MDPs can be extended to multi-agent settings, known as Markov games (MGs) (Littman, 1994) or stochastic games (Shapley, 1953). We focus on the most general form: partially observable Markov games (POMGs), where multiple agents interact with the environment, and each agent receives only individual observations instead of the full state, meaning they have limited access to the environment’s state.
@@
A POMG $\mathcal{G}$ is defined by $\mathcal{G}=(\mathcal{N},\mathcal{S},\mathcal{A},\mathcal{O},\mathcal{P},\mathcal{R},\gamma,\rho)$. $\mathcal{N}=\{1,\cdots,n\}$ denotes n agents. $\mathcal{S}$ is the state space. $\mathcal{A}=\prod_{i=1}^{n}\mathcal{A}_{i}$ is the product of the action space of each agent. Similarly, $\mathcal{O}=\prod_{i=1}^{n}\mathcal{O}_{i}$ is the product of the observation space of each agent. $\mathcal{P}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow[0,1]$ denotes the transition probability from one state to another given the actions of each agent. =$\mathcal{R}=\{\mathcal{R}_{1},\cdots,\mathcal{R}_{n}\}$, where $\mathcal{R}_{i}:\mathcal{S}\times\mathcal{A}_{i}\rightarrow\mathbb{R}$ denotes the reward function of agent $i$. $\gamma\in[0,1]$ is the discount factor. $\rho:\mathcal{S}\rightarrow[0,1]$ describes initial state distribution. If it is a fully cooperative setting, agents can share the same reward function (Yu et al., 2022; Rashid et al., 2020; Foerster et al., 2018; Kuba et al., 2021). When $n=1,\mathcal{O}_{i}=\mathcal{S}$, the POMG returns to the MDP.
@@
More concretely, in RL, agents interact with the environment through the following procedure: At each discrete time step $t$, each agent $i$ receives an observation $o_{i,t}$ from the environment and selects an action based on a stochastic policy $\pi_{\theta_{i}}:\mathcal{O}_{i}\times\mathcal{A}_{i}\rightarrow[0,1]$, where $\theta_{i}$ represents the parameters. After receiving the joint actions $\mathbf{a}_{t}=(a_{1,t},\cdots,a_{n,t})$, the environment transitions from the current state $s_{t}$ to a subsequent state $s_{t+1}$ according to the transition function $\mathcal{P}$ and sends a reward $r_{i,t+1}$ to every agent $i$. The ultimate goal of agent $i$ is to maximize the expected discounted cumulative rewards: $\mathbb{E}_{\pi_{\theta_{i}}}[\sum_{t=0}^{\infty}\gamma^{t}r_{i,t}]$.
2.2 Game Theory Concepts
2.2.1 Perfect Information and Imperfect Information
In a game with perfect information, only one player moves at a time. Each player comprehensively understands the current game state, the complete history of moves made, and the potential future outcomes. If these conditions are not met, the game is considered to have imperfect information (Lucchetti, 2011; Mycielski, 1992). An example of a game with perfect information is Go. In Go, both players have full access to the game state at all times, including the positions of all stones and the identity of the next player. In contrast, Texas Hold’em poker is a game with imperfect information. Players cannot see their opponents’ private cards, creating uncertainty and requiring decisions based on probabilistic reasoning, bluffing, and interpreting others’ actions.
2.2.2 Normal-Form and Extensive-Form
@@
The normal form and extensive form are two distinct representations of games in game theory. If a game $\mathcal{G}$ is represented in the normal-form, it can be expressed by $\mathcal{G}=(\mathcal{N},\mathbf{\Pi},\mathbf{u})$. $\mathcal{N}=\{1,2,\cdots,n\}$ denotes the players. The set $\mathbf{\Pi}=\Pi_{1}\times\cdots\times\Pi_{n}$ represents the pure strategy space for all players. A pure strategy specifies a particular and deterministic action for a player in the game. A vector $\bm{\pi}=(\pi_{1},\dots,\pi_{n})\in\mathbf{\Pi}$ is called a strategy profile. A mixed strategy assigns a probability distribution over the set of pure strategies. For player $i$, a mixed strategy is represented by $\sigma_{i}\in\Delta(\Pi_{i})$, where $\Delta$ denotes the probability simplex. $\mathbf{u}=(u_{1},\cdots,u_{n})$, where $u_{i}:\mathbf{\Pi}\rightarrow\mathbb{R}$, is a utility function that assigns a real-valued payoff to each player $i$. If $\forall\bm{\pi}\in\mathbf{\Pi},\sum_{i}u_{i}(\bm{\pi})=0$, the game is a zero-sum game, otherwise it is a general-sum game. If $\Pi_{1}=\cdots=\Pi_{n}$ and the payoffs are invariant under any permutation of the players’ strategies, the game is a symmetric game. If a finite set of players (especially two players) are involved and each player has a finite set of strategies, a normal-form game can be directly depicted in a matrix.
@@
In two-player zero-sum symmetric normal-form games, both players share the same pure strategy space, denoted by $\Pi$, such that $\Pi=\Pi_{1}=\Pi_{2}$. Since the utility function satisfies $u_{1}(\pi_{i},\pi_{j})=-u_{2}(\pi_{i},\pi_{j})$, we can simplify the utility to a single function $u$, where for $\pi_{i},\pi_{j}\in\Pi$, if $\pi_{i}$ beats $\pi_{j}$, then $u(\pi_{i},\pi_{j})=-u(\pi_{j},\pi_{i})>0$. The evaluation matrix captures the game outcomes by detailing the results of different strategies when they are played against each other: $A_{\Pi}=\{u(\pi_{i},\pi_{j}):\pi_{i},\pi_{j}\in\Pi\times\Pi\}$.
@@
If a game is represented in the extensive-form, it is expressed sequentially, illustrating the sequence of moves, choices made by the players, and the information available to each player during decision-making. Typically, a game in the extensive-form is represented by a game tree. This tree demonstrates the sequential and potentially conditional nature of decisions. Moreover, if player $i$ has perfect recall, it means that player $i$ remembers which action they have taken in the past. A game $\mathcal{G}$ represented in the extensive-form can be expressed by $\mathcal{G}=(\mathcal{N}\cup\{c\},H,Z,P,\mathcal{I},A,\mathbf{u})$. $\mathcal{N}=\{1,2,\cdots,n\}$ denotes a set of players. $c$ is chance node and can be regarded as a special player. $H$ represents a set of possible histories and $Z\subseteq H$ is a set of terminal histories. Order of moves is represented by a function $P(h)\in\mathcal{N}\cup\{c\}$ to indicate which player is to move, where $h\in H$. $\mathcal{I}$ denotes information set partitions and $\mathcal{I}_{i}$ denotes the information set partitions for player $i$. This implies that in an imperfect information game, if player $i$ reaches a history $h\in I_{i}$, where $I_{i}\in\mathcal{I}_{i}$ is a specific information set, player $i$ cannot distinguish which particular history $h\in I_{i}$ it is encountering. Action space is represented by $A(h)$ for a non-terminal history $h\in H$. For all non-terminal histories $h$ within an information set $I_{i}$, the available actions are the same; otherwise, they are distinguishable. Therefore, we use $A(I_{i})$ to represent the available actions for the information set $I_{i}$. Utility functions is denoted by $\mathbf{u}=(u_{1},\cdots,u_{n})$, where $u_{i}:Z\rightarrow\mathbb{R}$. Together, these components define the structure and dynamics of an extensive-form game. Moreover, A strategy profile in an extensive-form game can be expressed by $\bm{\pi}=(\pi_{1},\cdots,\pi_{n})$, where $\pi_{i}$ maps each $I_{i}\in\mathcal{I}_{i}$ to a probability distribution over $A(I_{i})$. A subgame of an extensive-form game is a portion of the game that starts from a single initial node, includes all successors of any node within the subgame, and contains all nodes in the same information set as any node in the subgame.
The Prisoner’s Dilemma is a classic example in game theory. The game outcomes are as follows:
  • If one player confesses (C) and the other lies (L), the confessor will serve 1 year in jail, while the liar will serve 8 years.
  • If both players choose to confess, they will each serve 7 years behind bars.
  • If both players choose to lie, they will each serve only 2 years behind bars.
The classic scenario is known as the simultaneous Prisoner’s Dilemma, where two players must simultaneously decide whether to confess or lie without knowing the other’s choice. Games played in this manner are referred to as static games. The normal-form representation is suitable for them, as it captures the simultaneous nature of decision-making (as depicted in Fig. 2a). Another variant is the sequential Prisoner’s Dilemma, where the second player decides with knowledge of the first player’s action. Games of this nature are called dynamic games. The extensive-form representation is well-suited for them, as it can clearly illustrate the sequence of moves and the information available to each player at each decision point (as shown in Fig. 2b).
Figure 2: The example of Prisoner’s Dilemma. (a) Matrix representation of simultaneous Prisoner’s Dilemma in normal-form. (b) Game tree representation of sequential Prisoner’s Dilemma in extensive-form.
Beyond normal-form and extensive-form games, the analysis of complex games often involves a higher-level abstraction: the meta-game. The meta-game can be viewed as an advanced form of a normal-form game, but it focuses on the selection of policies (also called strategies) rather than isolated actions. Each player chooses from a set of policies, rather than a set of actions. This set of policies is referred to as the policy population. Meta-strategies are mixed strategies that assign probabilities to the policies within the policy population in the meta-game.
2.2.3 Transitive Game and Non-transitive Game

### v4

@@ Abstract.
Self-play, a learning paradigm where agents iteratively refine their policies by interacting with historical or concurrent versions of themselves or other evolving agents, has shown remarkable success in solving complex non-cooperative multi-agent tasks. Despite its growing prominence in multi-agent reinforcement learning (MARL), such as Go, poker, and video games, a comprehensive and structured understanding of self-play remains lacking. This survey fills this gap by offering a comprehensive roadmap to the diverse landscape of self-play methods. We begin by introducing the necessary preliminaries, including the MARL framework and basic game theory concepts. Then, it provides a unified framework and classifies existing self-play algorithms within this framework. Moreover, the paper bridges the gap between the algorithms and their practical implications by illustrating the role of self-play in different non-cooperative scenarios. Finally, the survey highlights open challenges and future research directions in self-play.
Self-play, reinforcement learning, game theory
† Corresponding authors. This research was supported by National Natural Science Foundation of China (No.62406159, 62325405), Postdoctoral Fellowship Program of CPSF under Grant Number (GZC20240830, 2024M761676), China Postdoctoral Science Special Foundation 2024T170496.
copyright: noneccs: Computing methodologies Multi-agent systemsccs: General and reference Surveys and overviews
1. Introduction
Figure 1. Overview of our survey.
@@ 1. Introduction
Reinforcement learning (RL) represents a significant paradigm (Sutton and Barto, 2018) within machine learning (ML), focused on optimizing decision processes through interaction with an environment. In RL, the environment is modeled as a Markov decision process (MDP), where agents observe states, take actions, receive rewards, and cause transitions to new states. The primary goal of RL algorithms is to derive the optimal policy that yields the maximum expected accumulated reward over time. Deep RL extends traditional RL by employing deep neural networks as function approximators to handle high-dimensional state spaces, contributing to breakthroughs in various complex tasks (Mnih et al., 2013).
Transitioning from single-agent to multi-agent reinforcement learning (MARL) introduces complex dynamics (Rashid et al., 2018; Mahajan et al., 2019; Yu et al., 2022). In MARL, the interdependence of agents’ actions presents significant challenges, as the environment appears non-stationary to each agent. In the context of MARL, environments can generally be categorized into cooperative and non-cooperative settings. In cooperative scenarios, all agents share a common objective and are incentivized to work together, often leading to formulations that maximize a joint reward. In contrast, non-cooperative settings involve agents with individual or even conflicting goals, where the success of one agent may come at the expense of others. In MARL, non-cooperative tasks are generally harder than cooperative ones because the absence of a shared reward prevents the effective use of centralized training, a key technique for mitigating non-stationarity in cooperative settings (Foerster et al., 2016, 2018; Lowe et al., 2017; Yu et al., 2022), thereby further aggravating non-stationarity.
@@
With the help of game theory, a mathematical framework that models the interactions between multiple decision-makers, self-play emerges as an elegant solution to some inherent challenges in MARL. Vanilla self-play refers to a setting in which an agent continuously trains by competing against the most recent version of itself, as famously demonstrated for the symmetric zero-sum game of checkers (Samuel, 1959). Self-play, however, encompasses far more than this canonical example. More broadly, self-play refers to a class of learning methods where at least one agent iteratively improves its policy by interacting with a distribution of evolving opponents, which may include historical or concurrent versions of itself or other agents. It is worth noting that although some MARL studies have applied the concept of self-play in cooperative tasks such as Hanabi (Cui et al., 2021) and Overcooked (Strouse et al., 2021), our survey is conducted within non-cooperative settings. Self‑play has demonstrated remarkable success in non-cooperative domains such as Go (Silver et al., 2016, 2017, 2018; Schrittwieser et al., 2020), chess (Silver et al., 2018; Schrittwieser et al., 2020), poker (Moravčík et al., 2017; Heinrich et al., 2015), and complex video games (Berner et al., 2019; Vinyals et al., 2019). In these scenarios, it has developed strategies that surpass human expertise. Although the application of self-play is extensive and promising, it is accompanied by limitations, such as the potential convergence to suboptimal strategies and significant computational requirements (Silver et al., 2016, 2018).
@@
Although some research (Wellman et al., 2025) adopts a different perspective through empirical game-theoretic analysis (EGTA), relatively few comprehensive surveys focus on self-play. Among them, a study (Hernandez et al., 2021) develops an algorithmic framework for self-play, but it does not incorporate the Policy-Space Response Oracle (PSRO) (Lanctot et al., 2017) series of algorithms, while a separate study (Bighashdel et al., 2024) focuses exclusively on PSRO, without considering other self-play algorithms. Another study (DiGiovanni and Zell, 2021) examines the theoretical safety of self-play. While these studies are valuable, they do not provide a comprehensive perspective that fully captures the breadth and depth of self-play. To address this gap, this survey aims to offer a systematic roadmap to the diverse landscape of self-play algorithms in MARL within non-cooperative settings.
@@
The survey is organized as follows. Sec. 2 introduces the background of self-play, including the MARL framework and game theory concepts. Sec. 3 proposes a unified framework and then categorizes existing self-play algorithms into four categories based on this framework. In Sec. 4, a comprehensive analysis illustrates how self-play is applied in various scenarios. Sec. 5 describes open problems in self-play and explores future research directions. Finally, Sec. 6 concludes the survey on self-play. A more detailed overview of our survey is depicted in Fig. 1, which further illustrates the relationships among the subsections.
2. Preliminaries
@@
This section first introduces the framework of MARL, where agents interact with an environment to learn policies that maximize cumulative rewards. We then present key concepts from non-cooperative game theory, which provides tools to analyze strategic interactions among rational decision-makers with individual objectives.
2.1. MARL Framework
@@
In RL, the environment is modeled as an MDP, where the Markovian assumption states that the environment’s evolution is fully determined by its current state, eliminating the need to consider past states. An agent interacts with the environment by taking actions, which result in different states and associated rewards. MDPs can be extended to multi-agent settings, known as Markov games (MGs) (Littman, 1994) or stochastic games (Shapley, 1953). We focus on the most general form: partially observable Markov games (POMGs), where multiple agents interact with the environment, and each agent receives only individual observations instead of the full state, thus having limited access to the environment’s state.
@@
A POMG $\mathcal{G}$ is defined by $\mathcal{G}=(\mathcal{N},\mathcal{S},\mathcal{A},\mathcal{O},\mathcal{P},\mathcal{R},\gamma,\rho)$. $\mathcal{N}=\{1,\cdots,n\}$ denotes n agents. $\mathcal{S}$ is the state space. $\mathcal{A}=\prod_{i=1}^{n}\mathcal{A}_{i}$ is the product of the action space of each agent. Similarly, $\mathcal{O}=\prod_{i=1}^{n}\mathcal{O}_{i}$ is the product of the observation space of each agent. $\mathcal{P}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow[0,1]$ denotes the transition probability from one state to another given the actions of each agent. $\mathcal{R}=\{\mathcal{R}_{1},\cdots,\mathcal{R}_{n}\}$, where $\mathcal{R}_{i}:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ denotes the reward function of agent $i$. $\gamma\in[0,1]$ is the discount factor. $\rho:\mathcal{S}\rightarrow[0,1]$ describes initial state distribution.
@@
More concretely, in RL, agents interact with the environment through the following procedure: At each discrete time step $t$, each agent $i$ receives an observation $o_{i,t}$ from the environment and selects an action $a_{i,t}$ based on a stochastic policy $\pi_{i}:\mathcal{O}_{i}\times\mathcal{A}_{i}\rightarrow[0,1]$. After receiving the joint actions $\mathbf{a}_{t}=(a_{1,t},\cdots,a_{n,t})$, the environment transitions from the current state $s_{t}$ to a subsequent state $s_{t+1}$ according to the transition function $\mathcal{P}$ and sends a reward $r_{i,t+1}$ to every agent $i$. The ultimate goal of agent $i$ is to learn a policy $\pi_{i}$ that maximize the expected discounted cumulative rewards: $\mathbb{E}_{(\pi_{i},\pi_{-i})}[\sum_{t=0}^{\infty}\gamma^{t}r_{i,t+1}]$, where $\pi_{-i}$ denotes the other agents’ policies.
2.2. Game Theory Concepts
2.2.1. Normal-Form and Extensive-Form
@@
The normal form and extensive form are two distinct representations of games in game theory. If a game $\mathcal{G}$ is represented in the normal-form, it can be expressed by $\mathcal{G}=(\mathcal{N},\mathbf{\Pi},\mathbf{u})$. $\mathcal{N}=\{1,2,\cdots,n\}$ denotes the players. The set $\mathbf{\Pi}=\Pi_{1}\times\cdots\times\Pi_{n}$ represents the pure strategy space for all players. A pure strategy specifies a deterministic action choice for each point in the game. A vector $\boldsymbol{\pi}=(\pi_{1},\dots,\pi_{n})\in\mathbf{\Pi}$ is called a strategy profile. A mixed strategy $\boldsymbol{\sigma}$ assigns a probability distribution over the set of pure strategies. For player $i$, a mixed strategy is represented by $\sigma_{i}\in\Delta(\Pi_{i})$, where $\Delta$ denotes the probability simplex. $\mathbf{u}=(u_{1},\cdots,u_{n})$, where $u_{i}:\mathbf{\Pi}\rightarrow\mathbb{R}$, is a utility function that assigns the expected payoff to each player $i$. If $\exists C\in\mathbb{R},\forall\boldsymbol{\pi}\in\mathbf{\Pi},\sum_{i}u_{i}(\boldsymbol{\pi})=C$, the game is a constant-sum game. Specially, if $C=0$, it is a zero-sum games, in which every gain by one player is exactly offset by losses to the others. If no such constant $C$ exists, it is a general-sum game. Moreover, if $\Pi_{1}=\cdots=\Pi_{n}$ and for any permutation $\tau$ and any strategy profile $(\pi_{1},\dots,\pi_{n})$, $u_{i}(\pi_{1},\cdots,\pi_{n})=u_{\tau(i)}(\pi_{\tau(1)},\cdots,\pi_{\tau(n)})$, the game is a symmetric game. If a finite set of players is involved and each player has a finite set of strategies, a normal-form game can be depicted in a tensor $T\in\mathbb{R}^{|\Pi_{1}|\times\cdots\times|\Pi_{n}|\times n}$, where $|\Pi_{i}|$ is the size of player $i$’s strategy space and the final dimension indexes each player’s payoff.
@@
Specifically, in two-player zero-sum symmetric normal-form games, both players share the same pure strategy space, denoted by $\Pi$, such that $\Pi=\Pi_{1}=\Pi_{2}$. Since the utility function satisfies $u_{1}(\pi_{i},\pi_{j})=-u_{2}(\pi_{i},\pi_{j})$, we can simplify the utility to a single function $u$, where for $\pi_{i},\pi_{j}\in\Pi$, if $\pi_{i}$ beats $\pi_{j}$, then $u(\pi_{i},\pi_{j})=-u(\pi_{j},\pi_{i})>0$. This game can be directly depicted in an evaluation matrix $M_{\Pi}\in\mathbb{R}^{|\Pi|\times|\Pi|}$. It captures the game outcomes by detailing the results of different strategies when they are played against each other: $M_{\Pi}=\{u(\pi_{i},\pi_{j}):\pi_{i},\pi_{j}\in\Pi\times\Pi\}$.
@@
If a game is represented in the extensive-form, it is expressed sequentially, illustrating the sequence of moves, choices made by the players, and the information available to each player during decision-making. Typically, a game in the extensive-form is represented by a game tree. This tree demonstrates the sequential and potentially conditional nature of decisions. A game $\mathcal{G}$ represented in the extensive-form can be expressed by $\mathcal{G}=(\mathcal{N}\cup\{c\},H,Z,P,\mathcal{I},A,\mathbf{u})$. $\mathcal{N}=\{1,2,\cdots,n\}$ denotes a set of players. $c$ is the chance node and can be regarded as a special player. $H$ represents a set of possible histories and $Z\subseteq H$ is a set of terminal histories. Order of moves is represented by a function $P(h)\in\mathcal{N}\cup\{c\}$ to indicate which player is to move, where $h\in H$. $\mathcal{I}$ denotes information set partitions and $\mathcal{I}_{i}$ denotes the information set partitions for player $i$. If player $i$ reaches a history $h\in I_{i}$, where $I_{i}\in\mathcal{I}_{i}$ is a specific information set for player $i$, all decision nodes in $I_{i}$ are observationally equivalent to player $i$, and he cannot tell which specific node within $I_{i}$ has been reached. A game has perfect information precisely when every information set of every player is a singleton. Formally, $\forall i,\ \forall I_{i}\in\mathcal{I}_{i},\ |I_{i}|=1$. In such games, whenever a player is to move, he knows exactly which decision node has been reached and, equivalently, the full sequence of past actions and observed states. Otherwise the game is considered to have imperfect information. Action space is represented by $A(h)$ for a non-terminal history $h\in H$. For all non-terminal histories $h$ within an information set $I_{i}$, the available actions are the same; otherwise, they are distinguishable. Therefore, we use $A(I_{i})$ to represent the available actions for the information set $I_{i}$. Utility functions is denoted by $\mathbf{u}=(u_{1},\cdots,u_{n})$, where $u_{i}:Z\rightarrow\mathbb{R}$. Together, these components define the structure and dynamics of an extensive-form game. Moreover, a behavior strategy profile can be expressed by $\boldsymbol{\beta}=(\beta_{1},\cdots,\beta_{n})$, where $\beta_{i}$ maps each $I_{i}\in\mathcal{I}_{i}$ to a probability distribution over $A(I_{i})$. A game satisfies perfect recall if each player $i$ remembers the entire sequence of his past actions and all of the information he had at each of his previous decision points. In any finite extensive-form game with perfect recall, for every mixed strategy $\boldsymbol{\sigma}$, there exists an equivalent behavior strategy $\boldsymbol{\beta}$, and vice versa, such that both strategies induce the same distribution over terminal histories (i.e., the same outcomes in the game tree). A subgame of an extensive-form game is any subset of the game tree that (i) has exactly one initial decision node, (ii) whenever it contains a node it also contains all of its descendants, and (iii) whenever it contains a node it contains every other node in that node’s information set.
Beyond normal-form and extensive-form games, the analysis of complex games often involves a higher-level abstraction: the meta-game. A meta-game provides an abstraction over the original game by modeling interactions between entire policies rather than individual actions. While commonly instantiated as a normal-form game, a meta-game can, in general, take any game-theoretic form (e.g., extensive-form) as long as it captures strategic interactions over a policy population. In this setting, players choose policies instead of individual actions, and meta-strategies correspond to mixed strategies defined over these policies.
2.2.2. Transitive Game and Non-Transitive Game